OpenStudy (anonymous):

What are the differences among expressions, equations, and functions? Provide examples of each.

OpenStudy (anonymous):

expressions 2x - 11 \[\sqrt{412 - 2xy + 7}\] 45q

OpenStudy (anonymous):

equations 4x = 3y + 17 Area = pi x r^2 x^2 + y^2 = z^2

OpenStudy (anonymous):

functions f(x) = 2x-17 G(p) = 3p + 2p^2 - 17

OpenStudy (anonymous):

Thanks so much....that made it much easier for me to understand